OpenStudy (anonymous):

What values best approximate the solutions of x^2 + 4x - 3.5 = 0?

OpenStudy (anonymous):

help?

OpenStudy (anonymous):



OpenStudy (anonymous):

-4.8 and 3.5 ?

OpenStudy (anonymous):

\[x^2+4x-\frac{35}{10}\text{ = }0 \]The following are exact symbolic solutions to the above equation.\[\left\{x\to \frac{1}{2} \left(-4-\sqrt{30}\right),x\to \frac{1}{2} \left(-4+\sqrt{30}\right)\right\} \]\[\{x\to -4.73861,x\to 0.738613\} \]

OpenStudy (anonymous):

no real soultions?

OpenStudy (anonymous):

-4.8 and -3.5 -4.8 and 0.8 - 3.5 and 0.8

OpenStudy (anonymous):

@KierseyClemons If the above list represents three multiple choices for an answer, then the middle one is the best choice. -4.74 and 0.74 are accurate to the nearest 1/100th.

OpenStudy (anonymous):

ok! and also the other choice was no soloution